A 6-m-hight, 4-m-wide rectangular gate is located
Answer #1

Using equiibrium of momentum @ the hinge,

Fwater* (6-2) = F*(3)

Force required, F = Fwater*(4/3) = 4/3*Pavg * A=(4/3)*delta(P) /2* A=2/3* (P1-P0)*A